OpenStudy (anonymous):

1 over 3 (2x − 8) = 4

OpenStudy (anonymous):

are you trying to find the value of x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if that is the case you can invert both sides to make it\[3(2x-8)=\frac{ 1 }{ 4 }\] and solve from there

OpenStudy (anonymous):

thts confusing....

OpenStudy (goldphenoix):

I believe it's 1/3, not 3.

OpenStudy (goldphenoix):

(1/3)*

OpenStudy (anonymous):

wats the answer

OpenStudy (anonymous):

the original problem is this correct? \[\frac{ 1 }{ 3(2x-8) }=4\]

OpenStudy (goldphenoix):

Do the distributive property.

OpenStudy (anonymous):

that works too

OpenStudy (anonymous):

omg no help never mind

OpenStudy (goldphenoix):

\[\large \large \frac{ 1 }{ 3 }(2x-8)=4\]

OpenStudy (goldphenoix):

Answering is not helping. :|

OpenStudy (anonymous):

soo pleaseeeee

OpenStudy (anonymous):

im guessing this is for homework? alright so it should be \[(\frac{ 1 }{ 12 }+8)/2\]